# This work has been submitted to an IEEE conference for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. Mobile Jamming Mitigation in 5G Networks: A MUSIC-Based Adaptive Beamforming Approach

### v1

@@ Abstract
Mobile jammers pose a critical threat to 5G networks, particularly in military communications. We propose an intelligent anti-jamming framework that integrates Multiple Signal Classification (MUSIC) for high-resolution Direction-of-Arrival (DoA) estimation, Minimum Variance Distortionless Response (MVDR) beamforming for adaptive interference suppression, and machine learning (ML) to enhance DoA prediction for mobile jammers. Extensive simulations in a realistic highway scenario demonstrate that our hybrid approach achieves an average Signal-to-Noise Ratio (SNR) improvement of 9.58 dB (maximum 11.08 dB) and up to 99.8% DoA estimation accuracy. The framework’s computational efficiency and adaptability to dynamic jammer mobility patterns outperform conventional anti-jamming techniques, making it a robust solution for securing 5G communications in contested environments.
@@ I Introduction
Conventional interference mitigation strategies, such as fixed beamforming patterns, are effective against stationary jammers but fail to adapt to the mobility of jammers, which degrade network performance by introducing intentional interference across the spectrum [1]. Recent advances in adaptive beamforming, particularly combining Multiple Signal Classification (MUSIC) for Direction-of-Arrival (DoA) estimation with Minimum Variance Distortionless Response (MVDR) for interference suppression, offer promising solutions. However, these approaches often lack the real-time adaptability needed to counter unpredictable jammer mobility patterns, a critical gap for ensuring robust 5G communications in dynamic environments [2].
@@
This paper proposes an intelligent anti-jamming framework to address mobile jammers in 5G networks, leveraging adaptive beamforming optimized for dynamic scenarios, as shown in Fig. 1. Our approach integrates an enhanced MUSIC algorithm for high-resolution DoA estimation, MVDR beamforming for adaptive interference suppression, and machine learning (ML) to improve DoA prediction under varying jammer mobility. Key contributions include: (1) a novel MUSIC-ML hybrid for accurate DoA estimation, (2) tailored MVDR beamforming for mobile jammer mitigation, and (3) comprehensive performance analysis demonstrating significant Signal-to-Noise Ratio (SNR) improvements. The framework demonstrates robust performance in mobile jamming scenarios while maintaining computational efficiency.
The remainder of this paper is structured in the following way: Section II reviews background and related work. Section III describes the proposed MUSIC-MVDR-ML framework. Section IV presents experimental results, discussion, and limitations. Finally, Section V concludes the paper and outlines future research directions.
@@ II Background and Related Work
Beamforming, essential for enhancing signal quality and mitigating interference in modern wireless communications like 5G, has evolved from conventional fixed-weight methods to adaptive solutions [3]. While conventional beamforming offers simplicity, its fixed patterns struggle against dynamic interference, particularly mobile jammers. Adaptive techniques like MVDR and Linear-Constrained Minimum Variance (LCMV) adjust parameters based on the signal environment.
MVDR beamforming optimizes array weights to maximize sensitivity towards a desired DoA while suppressing other sources [4]. However, its performance degrades if the desired DoA is inaccurate. To address this, MVDR is coupled with high-resolution DoA estimation algorithms like MUSIC, which uses eigenspace decomposition to accurately identify signal directions even in multi-source scenarios [5, 6]. LCMV also minimizes interference but imposes linear constraints to preserve the target signal [7]. A practical limitation is that LCMV is often restricted to a uniform linear array (ULA), whereas MVDR offers more flexibility, adapting well to a uniform rectangular array (URA) used in this work [6].
Traditional anti-jamming strategies focusing on spatial filtering are effective against stationary jammers but fail when confronting mobile ones that dynamically change positions [8]. At the network layer, multipath communications and secret sharing schemes enhance resilience by distributing data across multiple paths to counter jamming attacks [9, 10]. The combination of MUSIC’s accurate DoA estimation with ML support and MVDR’s adaptive nulling presents a promising approach for mobile interference mitigation.
Finally, accurately evaluating performance through signal-to-noise ratio (SNR) in multi-antenna systems facing mobile jammers presents challenges. Standard SNR calculations may not adequately handle dynamically changing interference or the specifics of signal combination after beamforming, necessitating careful consideration of the calculation methodology, as addressed in our proposed framework. The limitations of current DoA estimation with mobile sources, real-time adaptation challenges, and the need for better ML integration highlight the necessity for the comprehensive approach developed in this paper.
@@ III Proposed Framework
Our proposed framework integrates MUSIC-based DoA estimation, MVDR beamforming, and ML enhancement for DoA estimation support. We consider a $2\times 2$ URA that receives signals from a desired transmitter and a mobile jammer. The received signal $x(t)$ at the array is modeled as:
@@
where $s(t)$ is the desired signal, $j(t)$ is the jamming signal, $n(t)$ is additive white Gaussian noise, and $a(\theta,\phi)$ is the array steering vector which corresponds to azimuth angle $\theta$ and elevation angle $\phi$. Subscripts $s$ and $j$ denote the desired signal and jammer, respectively.
@@ III-A Anti-Jamming System
MUSIC-Enhanced DoA Estimation: The MUSIC algorithm estimates an angle of arrival by exploiting the eigenstructure of the received signal’s covariance matrix $R$ [5]:
@@
Eigendecomposition of $R$ separates the signal subspace ($U_{s}$) from the noise subspace ($U_{n}$) [11]:
@@
This allows for high-resolution DoA estimation by using $R$ and scanning over a grid of possible angles, in our case -90 to 90, then evaluating the MUSIC pseudospectrum. Peaks in the pseudospectrum show the estimated DoA of an incoming signal.
Adaptive MVDR Beamforming: Beamforming is a spatial filtering strategy where various weight geometries and gain tapering methods manipulate transmitter and receiver antennas. A conventional beamformer is designed to stay the same between collections, while an adaptive beamformer responds to environment, allowing the antenna array to dynamically change. The MVDR beamformer is an adaptive beamformer that computes optimal weights $w_{opt}$ to minimize output power while preserving the desired signal gain (in the direction $\theta_{s}$, assumed to be the desired signal’s DoA) [12]:
$w_{opt}=\frac{R^{-1}a(\theta_{s},\phi_{s})}{a^{H}(\theta_{s},\phi_{s})R^{-1}a(\theta_{s},\phi_{s})}$ (4)
This process adaptively creates nulls in the direction of interference sources such as the mobile jammer. In order to effectively shape the beam, we apply $w_{opt}$ to the antenna array.
Machine Learning Integration: To enhance estimation accuracy, particularly when the initial MUSIC DoA estimation error exceeds a predefined threshold, a linear regression ML model is employed. Linear regression was chosen due to its simplicity in implementation and relationship between input output. The model is trained using MATLAB’s fitlm [13] where the input features (predictors) are sequences of three consecutive ground truth azimuth angle values, and the output (response) is the subsequent ground truth azimuth value. Over 500 DoA data points from the simulated mobility patterns of the scenario were used as training data shown in Fig. 2. The model predicts the DoA’s next azimuth angle $\hat{\theta}_{t+1}$ based on the past $k$ measurements $(\theta_{t},\theta_{t-1},...,\theta_{t-k})$:
@@
While $k=3$ is only used in our experiments, the above linear prediction model can be easily generalized by leverging more historical measurements. The high statistical significance of the $\theta_{t-2}$ coefficient (p-value $\approx 3.4\times 10^{-76}$) in our experiments confirms strong temporal dependency, justifying the model’s predictive capability.
Figure 2: Optimized linear regression model showing relationship between the adjusted whole model values and the adjusted data with 95% confidence bounds.
@@ III-B Performance Quantification
SNR Calculation: To reliably evaluate performance, we use a specific SNR calculation method that averages antenna signals across the $M\times N$ array elements to prevent artificial gain inflation and ensures consistency between simulation and potential implementation, both before and after MVDR processing. The combined signal is $x_{2}(t)=\frac{1}{N+M}\sum_{i=1}^{N+M}|x(t)_{i}|$. Signal and noise powers are calculated by averaging $|x_{2}(t)|$ over known signal-present ($a_{1}$ to $b_{1}$) and signal-absent ($a_{0}$ to $b_{0}$) regions:
@@
The SNR improvement is calculated as:
$\text{SNR}_{improvement}=\text{SNR}_{after}-\text{SNR}_{before}$ (9)
Figure 3: Simulation framework for mobile jamming mitigation with ML-enabled dynamic beamforming.
Mitigation Algorithm: Fig. 3 is a flow diagram of Algorithm 1 which outlines the overall simulation framework to mitigate jamming in mobile environment, which integrates MUSIC-based DoA estimation, MVDR beamforming, and ML prediction model. It iteratively collects signals, estimates DoA using MUSIC (potentially refined by the ML model when MUSIC error is high, as described above), computes MVDR weights, applies beamforming, and evaluates performance based on the SNR improvement. The system can adapt parameters or retrain the ML model, if necessary, to further improve the performance. This integrated approach provides robust protection against mobile jammers while maintaining computational efficiency.
@@
8:     if $DoA$ $Error>\epsilon$ then
@@
11:     Update ML model with latest measurements (e.g., recent DoAs for next prediction)
@@
13:     Apply beamforming with $w_{\mathrm{opt}}$ and measure $\text{SNR}_{{after}}$ using Eq. (8)
14:     Calculate $\text{SNR}_{improvement}$ using Eq. (9)
@@ IV-A Setup
We evaluated the proposed framework through extensive MATLAB simulations emulating a realistic highway scenario as seen in Fig. 1. This scenario involved a stationary transmitter (cell tower), a mobile receiver (red car), and a mobile jammer (blue van) moving along various predefined paths. The receiver and jammer have the ability to begin in any of the three cardinal directions (where the receiver in the figure began on the left-hand side of the scenario, which is equivalent to west). Trajectory options include North-to-South, North-to-West, South-to-North, South-to-West, West-to-North, and West-to-South. Key simulation parameters include:
The carrier frequency of 100 MHz was selected to simplify simulation and reduce computational complexity, as the MUSIC-MVDR-ML framework is frequency-agnostic and can be adapted to standard 5G bands like 3.5 GHz in future implementations. Similarly, the signal sampling rate of 1 kHz was chosen for baseband signal processing after downconversion, which is sufficient for the simulated scenario.
@@ IV-B Results
DoA Estimation and ML Performance: The integrated MUSIC algorithm demonstrated high DoA estimation accuracy (up to 99.8%) with a dynamic jammer present across the tested scenarios. Fig. 4 shows an example of a beam pattern with the estimated DoA indicated by its peak at -40. The machine learning support aids in this estimation when error is too high.
Adaptive Beamforming: The adaptive MVDR beamformer effectively steered nulls towards the jammer direction while preserving the desired signal. Analysis of the beam patterns (e.g., Fig. 4) confirmed improved directionality and interference suppression compared to initial states. The system successfully mitigated jamming interference along both linear and diagonal trajectories within the highway environment.
SNR Enhancement: Using our defined SNR calculation method, which averages the combined antenna signals, the framework achieved significant improvement. Across the six diverse highway trajectories simulated, the system yielded an average SNR improvement of 9.58 dB post-MVDR beamforming. While the improvement varied depending on the relative positions and path (as illustrated in Fig. 5), for all trajectories there is a spike in SNR at around sample 25 that corresponds to the intersection point of the simulated highway. The enhancement was consistently observed, validating the system’s anti-jamming capability.
@@
Computational Efficiency: The framework demonstrated suitability for real-time application, with average processing times per collection measured as: DoA estimation (MUSIC) $\approx$ 41.9 ms, MVDR weight computation $\approx$ 5.2 ms, and ML prediction $\approx$ 16.9 ms111github.com/CLIS-WPI/mobile-jamming-mitigation..
Baseline Comparison: As previously mentioned, adaptive beamforming is used in this anti-jamming system as well as machine learning. The following points highlight direct comparisons among different MATLAB simulations. The values indicate SNR improvement, DoA accuracy, and processing time across three methods:
  • MUSIC-MVDR-ML (Proposed): 9.58 dB, 99.8%, 63.0 ms
  • MUSIC-MVDR: 9.46 dB, 99.3%, 74.4 ms
  • Fixed Beamforming: 1.90 dB, 88.0%, 240.2 ms
@@ IV-C Discussion
Limitations: Despite robust performance, limitations exist in this adaptive beamforming for anti-jamming setup. SNR calculation accuracy can decrease in very high-mobility scenarios. The current ML model might require retraining for drastically different mobility patterns not encountered during training. Furthermore, the framework’s performance against multiple simultaneous jammers needs investigation. Only one jammer was implemented for sake of simplicity. To introduce a multi-jammer system more jammer objects and channels would need to be initalized and added at collection time. Lastly, due to the nature of the simulation, the propagation of signals occurs at every collection prior to the processing step inside the iterative for loop. This causes latency within the system, as multi-threading was not explored.
@@
TABLE I: Summary of Key Performance Metrics
@@
V Conclusion and Future Work
This paper presented a novel anti-jamming framework integrating MUSIC DoA estimation, MVDR beamforming, and machine learning to mitigate mobile 5G jammers. Simulations confirmed its effectiveness, achieving an average 9.58 dB SNR improvement and high DoA accuracy (up to 99.8%) in realistic highway scenarios. Future work will focus on addressing the limitations, potentially exploring deep learning models and optimizing for multi-jammer environments. This project demonstrated adaptive beamforming for anti-jamming as a proof of concept in simulation. Next steps would be implementation in a lab setting with USRP software defined radios and PCB directional antennas.

### v2

# A MUSIC-Based Adaptive Beamforming Approach for Mobile Jamming Mitigation in 5G Networks

@@ Abstract
Mobile jammers pose a critical threat to 5G networks, particularly in military communications. This paper investigates an anti-jamming framework that enhances a strong adaptive beamforming baseline comprising Multiple Signal Classification (MUSIC) for Direction-of-Arrival (DoA) estimation and Minimum Variance Distortionless Response (MVDR) for interference suppression with a lightweight machine learning (ML) model for predictive error correction. Extensive simulations in a realistic highway scenario demonstrate that the integrated system achieves a high DoA estimation accuracy of up to 99.8% and an average Signal-to-Noise Ratio (SNR) improvement of 9.58 dB. Analysis reveals that the MUSIC-MVDR baseline alone accounts for the vast majority of this performance gain (9.46 dB), indicating that the primary benefit of the simple ML model lies in correcting outlier estimates rather than providing a substantial systemic SNR increase. The framework’s computational efficiency validates the effectiveness of the core beamforming approach and highlights the critical trade-off between ML model complexity and practical performance gains for securing 5G communications in contested environments.
@@ I Introduction
Conventional interference mitigation strategies, such as fixed beamforming patterns, are effective against stationary jammers but fail to adapt to the mobility of jammers, which degrade network performance by introducing intentional interference across the spectrum [9733393]. The combination of Multiple Signal Classification (MUSIC) for Direction-of-Arrival (DoA) estimation and Minimum Variance Distortionless Response (MVDR) for interference suppression is a well-established and powerful approach for adaptive beamforming. However, the performance of this combination can still be challenged by highly dynamic scenarios where DoA estimation errors may occur. This presents an opportunity to investigate whether lightweight computational methods can enhance the robustness of this framework without incurring significant processing overhead [10149863].
@@
This paper explores an enhanced anti-jamming framework designed for mobile jammers in 5G networks, as depicted in Fig. 1. The framework builds upon the MUSIC-MVDR baseline and integrates a lightweight machine learning (ML) model to improve DoA estimation accuracy, particularly in cases where the primary MUSIC algorithm yields high error. The key objectives are: (1) to implement a robust MUSIC-MVDR baseline for mobile jammer mitigation, (2) to assess the potential of a simple linear regression model to correct DoA estimation errors with minimal computational cost, and (3) to provide a comprehensive performance analysis quantifying the trade-offs between performance gain and complexity. The results demonstrate robust performance in mobile jamming scenarios while maintaining computational efficiency suitable for real-time applications.
The remainder of the paper is structured as follows: Section II reviews background and related work. Section III describes the proposed MUSIC-MVDR-ML framework. Section IV presents experimental results, discussion, and limitations. Finally, Section V concludes the paper and outlines future research directions.
@@ II Background and Related Work
Beamforming, essential for enhancing signal quality and mitigating interference in modern wireless communications like 5G, has evolved from conventional fixed-weight methods to adaptive solutions [benesty_brief_2021]. While conventional beamforming offers simplicity, its fixed patterns struggle against dynamic interference, particularly mobile jammers. Adaptive techniques like MVDR and Linear-Constrained Minimum Variance (LCMV) adjust parameters based on the signal environment.
MVDR beamforming optimizes array weights to maximize sensitivity towards a desired DoA while suppressing other sources [darzi_null_2014]. However, its performance degrades if the desired DoA is inaccurate. To address this, MVDR is commonly coupled with high-resolution DoA estimation algorithms like MUSIC, which uses eigenspace decomposition to accurately identify signal directions even in multi-source scenarios [1143830, 10366174]. LCMV also minimizes interference but imposes linear constraints to preserve the target signal [7471643]. A practical limitation is that LCMV is often restricted to a uniform linear array (ULA), whereas MVDR offers more flexibility, adapting well to the uniform rectangular array (URA) used in this study, making it a suitable choice for investigation [10366174].
Traditional anti-jamming strategies focusing on spatial filtering are effective against stationary jammers but fail when confronting mobile ones that dynamically change positions [10811542]. Recent research has explored integrating machine learning to address this limitation. While complex models like deep neural networks are being investigated, this study emphasizes the trade-off between performance and efficiency by employing a simple, interpretable ML model. At the network layer, multipath communications and secret sharing schemes also enhance resilience by distributing data across multiple paths to counter jamming attacks [10773801, 10773704]. The proposed approach complements these strategies by focusing on the physical layer, where the combination of MUSIC’s accurate DoA estimation, MVDR’s adaptive nulling, and supplementary ML support presents a practical method for mobile interference mitigation.
Finally, accurately evaluating performance through signal-to-noise ratio (SNR) in multi-antenna systems facing mobile jammers presents challenges. Standard SNR calculations may not adequately handle dynamically changing interference or the specifics of signal combination after beamforming, necessitating careful consideration of the calculation methodology, as addressed in the proposed framework. Limitations of current DoA estimation with mobile sources, real-time adaptation challenges, and the need for better ML integration highlight the necessity for the comprehensive approach developed in this study.
@@ III Proposed Framework
The proposed framework integrates MUSIC-based DoA estimation, MVDR beamforming, and ML enhancement for DoA estimation support. A $2\times 2$ URA receives signals from a desired transmitter and a mobile jammer. The received signal $x(t)$ at the array is modeled as:
@@
where $s(t)$ is the desired signal, $j(t)$ is the jamming signal, $n(t)$ is additive white Gaussian noise, and $a(\theta,\phi)$ is the array steering vector corresponding to azimuth angle $\theta$ and elevation angle $\phi$. Subscripts $s$ and $j$ denote the desired signal and jammer, respectively.
@@ III-A Anti-Jamming System
MUSIC-Enhanced DoA Estimation: The MUSIC algorithm estimates angles of arrival by exploiting the eigenstructure of the received signal’s covariance matrix $R$ [1143830]:
@@
Eigendecomposition of $R$ separates the signal subspace ($U_{s}$) from the noise subspace ($U_{n}$) [matlab_music]:
@@
High-resolution DoA estimation is achieved by scanning over a grid of possible angles, in this study from -90 to 90, and evaluating the MUSIC pseudospectrum. Peaks in the pseudospectrum indicate the estimated DoA of incoming signals.
Adaptive MVDR Beamforming: Beamforming is a spatial filtering strategy where weight geometries and gain tapering methods manipulate transmitter and receiver antennas. A conventional beamformer remains static, while an adaptive beamformer responds to environmental changes, allowing the antenna array to adjust dynamically. The MVDR beamformer computes optimal weights $w_{\text{opt}}$ to minimize output power while preserving the desired signal gain in the direction $\theta_{s}$ [matlab_mvdr]:
$w_{\text{opt}}=\frac{R^{-1}a(\theta_{s},\phi_{s})}{a^{H}(\theta_{s},\phi_{s})R^{-1}a(\theta_{s},\phi_{s})}$ (4)
This adaptively creates nulls in the direction of interference sources, such as the mobile jammer. The weights $w_{\text{opt}}$ are then applied to the antenna array for beam shaping.
Machine Learning Integration: To enhance DoA estimation accuracy when initial MUSIC estimates exceed a predefined error threshold, a linear regression ML model is employed. Linear regression is chosen for its simplicity and interpretability. The model is trained using MATLAB’s fitlm [matlab_ml], with input features as sequences of three consecutive ground truth azimuth angles and the output as the subsequent ground truth azimuth angle. Over 500 DoA data points from simulated mobility patterns are used as training data (Fig. 2). The model predicts the next azimuth angle $\hat{\theta}_{t+1}$ based on the past $k$ measurements $(\theta_{t},\theta_{t-1},...,\theta_{t-k})$:
@@
While $k=3$ is used in these experiments, the model can be generalized to include more historical measurements. The statistical significance of the $\theta_{t-2}$ coefficient (p-value $\approx 3.4\times 10^{-76}$) confirms strong temporal dependency, supporting the predictive capability of the model.
Figure 2: Optimized linear regression model showing the relationship between predicted and actual azimuth angles with 95% confidence bounds.
@@ III-B Performance Quantification
SNR Calculation: Performance evaluation uses an SNR calculation method that averages antenna signals across the $M\times N$ array elements to prevent artificial gain inflation and ensure consistency between simulation and potential implementation. The combined signal is $x_{2}(t)=\frac{1}{N+M}\sum_{i=1}^{N+M}|x(t)_{i}|$, and signal and noise powers are calculated over signal-present ($a_{1}$ to $b_{1}$) and signal-absent ($a_{0}$ to $b_{0}$) regions:
@@
SNR improvement is computed as:
$\text{SNR}_{\text{improvement}}=\text{SNR}_{\text{after}}-\text{SNR}_{\text{before}}$ (9)
Figure 3: Block diagram of the proposed MUSIC-MVDR-ML framework showing the step-by-step mitigation process.
Mitigation Algorithm: Fig. 3 illustrates Algorithm 1, which outlines the simulation framework for mitigating mobile jamming by integrating MUSIC-based DoA estimation, MVDR beamforming, and the ML prediction model. The algorithm iteratively collects signals, estimates DoA using MUSIC (optionally refined by the ML model when errors exceed the threshold), computes MVDR weights, applies beamforming, and evaluates performance based on SNR improvement. The system can adjust parameters or retrain the ML model if necessary to improve performance. This integrated approach provides robust protection against mobile jammers while maintaining computational efficiency.
@@
8:  if $\text{DoA Error}>\epsilon$ then
@@
11:  Update ML model with latest measurements for next prediction
@@
13:  Apply beamforming with $w_{\mathrm{opt}}$ and measure $\text{SNR}_{\text{after}}$
14:  Calculate $\text{SNR}_{\text{improvement}}$
@@ IV-A Setup
The proposed framework was evaluated through extensive MATLAB simulations emulating a realistic highway scenario (Fig. 1). The scenario involved a stationary transmitter (cell tower), a mobile receiver (red car), and a mobile jammer (blue van) moving along predefined paths. The receiver and jammer could begin from any of the three cardinal directions (in the figure, the receiver began on the west side). Trajectory options included North-to-South, North-to-West, South-to-North, South-to-West, West-to-North, and West-to-South. Key simulation parameters are as follows:
The carrier frequency of 100 MHz was selected to simplify the simulation and reduce computational complexity. The MUSIC-MVDR-ML framework is frequency-agnostic and can be adapted to standard 5G bands, such as 3.5 GHz, in future implementations. A signal sampling rate of 1 kHz was used for baseband processing after downconversion, sufficient for the simulated scenario.
@@ IV-B Results
DoA Estimation and ML Performance: The integrated MUSIC algorithm demonstrated high DoA estimation accuracy (up to 99.8%) in the presence of a dynamic jammer across the tested scenarios. This accuracy was achieved using a hierarchical error threshold system with tolerance levels of 15%, 10%, and 7.5% for combined azimuth/elevation estimation errors, calculated as the percentage deviation from the total angle span. The ML model activates when MUSIC fails three consecutive estimation attempts, using linear regression based on the previous three DoA measurements.
Fig. 4 shows an example beam pattern with the estimated DoA indicated by its peak at -40. Machine learning support aids the estimation when error exceeds a threshold.
Adaptive Beamforming: The adaptive MVDR beamformer effectively steered nulls towards the jammer direction while preserving the desired signal. Analysis of beam patterns (e.g., Fig. 4) confirmed improved directionality and interference suppression compared to initial states. The system mitigated jamming interference along both linear and diagonal highway trajectories.
SNR Enhancement: Using the defined SNR calculation method, which averages combined antenna signals, the framework achieved significant improvement. Across six diverse highway trajectories, the system yielded an average SNR improvement of 9.58 dB post-MVDR beamforming. While the improvement varied depending on relative positions and paths (Fig. 5), all trajectories exhibited a spike in SNR around sample 25, corresponding to the highway intersection. This enhancement demonstrates the framework’s anti-jamming capability.
@@
Computational Efficiency: The framework is suitable for real-time applications, with average processing times per collection as follows: DoA estimation (MUSIC) $\approx$ 41.9 ms, MVDR weight computation $\approx$ 5.2 ms, and ML prediction $\approx$ 16.9 ms111github.com/CLIS-WPI/mobile-jamming-mitigation..
Baseline Comparison: Adaptive beamforming and machine learning were compared across MATLAB simulations. Performance metrics for SNR improvement, DoA accuracy, and processing time for three methods are summarized in Table I.
TABLE I: Detailed Algorithm Comparison
Algorithm SNR (dB) Accuracy (%) Type
Fixed Beamforming 1.90 88.0 Static
MUSIC-MVDR 9.46 99.3 Adaptive
MUSIC-MVDR-ML 9.58 99.8 ML-Enhanced
@@ IV-C Discussion
The experimental results demonstrate the effectiveness of the adaptive beamforming framework. A critical analysis of components and limitations provides further insights.
Analysis of ML Component Impact: The baseline comparison shows that the MUSIC-MVDR framework alone achieves an average SNR improvement of 9.46 dB, while adding the ML model increases this to 9.58 dB. This marginal gain of 0.12 dB indicates that, in the simulated highway trajectories, the MUSIC algorithm is highly robust, leaving limited room for improvement by a simple predictive model.
The superiority of MUSIC-MVDR-ML stems from three key factors: (1) MUSIC provides sub-degree DoA resolution through eigenspace decomposition, (2) MVDR creates precise adaptive nulls using optimal weight computation $w_{\text{opt}}=R^{-1}a(\theta_{s},\phi_{s})/[a^{H}(\theta_{s},\phi_{s})R^{-1}a(\theta_{s},\phi_{s})]$, and (3) the ML model serves as an outlier correction mechanism when MUSIC estimation quality degrades. The modest ML improvement validates that the MUSIC-MVDR baseline already achieves near-optimal performance in single-jammer scenarios.
The linear regression model primarily acts as a corrective mechanism for outlier DoA estimates, improving accuracy from 99.3% to 99.8% while exerting minimal impact on SNR. This highlights a crucial trade-off between pursuing higher performance with complex ML models and maintaining a computationally efficient framework that ensures baseline effectiveness and reliability.
Limitations: Despite robust performance, several limitations remain. The relatively modest impact of the current ML model suggests that future work should explore more sophisticated architectures, such as Recurrent Neural Networks (RNNs) or Long Short-Term Memory (LSTM) networks, which can better capture temporal dependencies in jammer mobility patterns and potentially provide greater gains. Such models, however, require larger datasets and rigorous validation techniques (e.g., cross-validation) to prevent overfitting.
Additionally, performance against multiple simultaneous jammers has not yet been investigated, as only a single jammer was simulated. Multi-jammer scenarios would require modeling multiple jammer objects and channels in each collection. Finally, the serial simulation loop, in which signal propagation occurs at every collection before processing, introduces latency. Future implementations should consider multi-threading or parallel processing to achieve true real-time operation.
The controlled simulation environment with predictable highway mobility patterns also contributed to the high accuracy achieved. Real-world implementations may face additional challenges, including multipath propagation, non-Gaussian noise, and more complex jammer movement patterns that may affect the reported performance metrics.
@@
TABLE II: Summary of Key Performance Metrics
@@
V Conclusion
An adaptive beamforming framework combining MUSIC and MVDR was presented and evaluated for mitigating mobile jammers in 5G networks. The integration of a lightweight machine learning model was examined to supplement the core algorithm. The key finding indicates that, while the MUSIC-MVDR baseline provides strong and effective anti-jamming performance, the contribution of a simple linear regression model to overall SNR improvement is marginal. This suggests that significant enhancements in highly dynamic environments require more sophisticated predictive models. Future work will focus on implementing the framework on a hardware testbed using SDR and exploring advanced deep learning architectures to address multi-jammer scenarios and more complex mobility patterns.